2. The equations of motion of this system are; 12Y+7Y+247-62-122=0 62 +62 +12Z- 6Y-12Y=f(t) 7W+7W+14W-7Y-14Y=ft) Where
Answer #1

State Variable form: Given equation of motions are; 12Y + 7Y + 24Y – 6Z – 12Z = 0 6Z + 6Z + 12Z - 6Y – 12Y = f(t) 7W +

1 X2 = Z = 6 (-6Z – 12Z + 6Y + 12Y + f(t)) f(t) 6 = (-X2 – 2x3 + x2 + 2X. + 109) 1 X3 = W = (-7W – 14W+7Y + 14Y + f(t))

0 1 0 IrX1 1/6 X2 X2 X3 | 1/7 [f(t)] = ܢܢ + X4 The state space model is represented as, • Å = AX + Bu [X 7 1 0 -2 X2 12 2 1 -

• Y = CX + Du Y2 Xi X2 X3 Y3 = = [1 1 1 0 0 0] YA + [0]u Y LY LX -